Suppose that a Demand Curve is given by QY = 10 – 5PY + PX Where QY is the quantity of the good, PY is the price of the good and PX pertains to the price of a substitutable good. The price of the substitutable good is K2. Suppose the price of the good is K1. Find the own price elasticity of demand? Find the cross-price elasticity of demand?


1
Expert's answer
2021-11-21T16:52:47-0500

Given,

Demand curve of good Y:

"QY=10\u22125P_Y+P_X"

Where,

QY=Quantity of good Y

PY=Price of good Y

PX=Price of good X


PY=price of the good Y=K1

PX=price of the substitute good=K2


The following formulas will be used:

Price elasticity of demand "=\frac{∂Q_Y}{∂P_Y}×\frac{P_Y}{Q_Y}"

Cross−price elasticity of demand "=\frac{∂Q_Y}{∂P_x}×\frac{P_x}{Q_Y}"

Calculation of price elasticity of demand:

"\frac{∂Q_Y}{∂P_Y}=−5"

Let us substitute the values of PY and PX in the demand function for quantity

 of good Y

"Q_Y=10−5×1+2\\Q_Y=10−5+2\\Q_Y=7"

Price elasticity of demand"=-5\times{1}{7}"

Price elasticity of demand=−0.71

Calculation of cross-price elasticity of demand:

"\frac{∂Q_Y}{∂P_X}=1\\P_X=K2\\Q_Y=7"

Cross−price elasticity of demand"=1\times \frac{2}{7}"

Cross−price elasticity of demand=0.29